OpenStudy (kittykr98):

use the function y+2=-(x-3)^2 to answer the following. A) show or explain how you would find the vertex. B) show or explain how to find the axis of symmetry. C) show or explain how you would find the x-intercepts. D) show or explain how you would find the y-intercept. E) Graph the function.

OpenStudy (kittykr98):

@shelbygt520 @studygurl14 @jim_thompson5910 @jaeuni

OpenStudy (anonymous):

You need to set up the equations as follows y=(x-h)^2+k

OpenStudy (kittykr98):

which would be converting the function to standard form, correct?

OpenStudy (anonymous):

Vertex form...

OpenStudy (kittykr98):

to do that, i would have to subtract 2 from both sides. giving me y=-(x-3)^2 -2 right?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

(h,k) is your vertex

OpenStudy (kittykr98):

so 3 and 2? or is it -3?

OpenStudy (kittykr98):

-3 and -2

OpenStudy (anonymous):

It is (3,-2)

OpenStudy (kittykr98):

ok so for A, i need to write "To find the vertex of the function, the function must first be converted to vertex form, y=(x-h)^2+k. To do this, we must subtract 2 from both sides of the function. This gives us y=-(x-3)^2-2."

OpenStudy (kittykr98):

I have to show work

OpenStudy (anonymous):

Axis of symmetry is k

OpenStudy (kittykr98):

im still kinda hung up on the vertex part. The vertex is (3, -2) but how do i list my work for that?

OpenStudy (anonymous):

that looks good

OpenStudy (kittykr98):

oh ok

OpenStudy (kittykr98):

ok the line axis of symmetry.. the equation is x= -b/2a right?

OpenStudy (kittykr98):

what is the formula for how to get the vertex?

OpenStudy (anonymous):

Axis of symmetry is h sorry my error.

OpenStudy (anonymous):

or x=h

OpenStudy (kittykr98):

I really think i need to show more work for how i got the vertex. they asked us to show work

OpenStudy (anonymous):

y=(x-h)^2+k this is the formula in vertex form

OpenStudy (kittykr98):

Yeah and i converted the function the vertex form. But how do i know which is the vertex?

OpenStudy (anonymous):

All you have to do is show the values that correspond to h and k. The vertex is x= h

OpenStudy (kittykr98):

OK how do i show the values that correspond?

OpenStudy (anonymous):

By the formula

OpenStudy (kittykr98):

ok i did this last night but i found the vertex by substituting the value of the line of symmetry into the equation but this time it doesnt ask for the line of symmetry until after finding the vertex

OpenStudy (kittykr98):

so thats why im having trouble with this

OpenStudy (anonymous):

y intercept is (0,k)

OpenStudy (anonymous):

for the x intercept make y = 0 and solve for x using the quadratic formula

OpenStudy (anonymous):

Last night you were using the standard form probably..this is the vertex form

OpenStudy (kittykr98):

you're way ahead of me..

OpenStudy (anonymous):

y=(x-h)^2+k vertex form y = ax^2+bx+c standard form

OpenStudy (kittykr98):

i got that part

OpenStudy (anonymous):

So when you subtracted 2 from both sides you ended up with this y=-(x-3)^2-2 correct?

OpenStudy (kittykr98):

yes

OpenStudy (anonymous):

So we are told that (h,k) represent the vertex. The axis of symmetry is x = h The y intercept is (0,k)

OpenStudy (kittykr98):

so the axis of symmetry is h which is 3 in this case?

OpenStudy (kittykr98):

and the y intercept is -2

OpenStudy (anonymous):

(0,-2)

OpenStudy (anonymous):

now take y=-(x-3)^2-2 and make y = 0 and solve to find your x intercepts

OpenStudy (kittykr98):

wait so make x=0?

OpenStudy (kittykr98):

im confused

OpenStudy (anonymous):

To find x intercepts you make y = 0

OpenStudy (kittykr98):

How would i do that????

OpenStudy (anonymous):